# Distance-based measures and Epsilon-measures for measurement-based quantum resources

### v3

@@ III.2 Distance-based measures for measurement-based resources
For the simplicity of notation in the following arguments let us denote $\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}}$ as $\mathcal{M}^{\prime}$. Then, we have ^$\hat{\mathcal{M}^{\prime}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{C}}=\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{C}}$ and therefore we can write the following,
@@
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})\geq$ $\mathbbm{K}(\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}},\mathcal{H}_{C})\leavevmode\nobreak\ \forall\mathcal{H}_{C}$
$or,\leavevmode\nobreak\ \inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})\geq$ $\inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}},\mathcal{H}_{C})$
$or,\leavevmode\nobreak\ \overline{\mathbbm{R}}(\mathcal{M})\geq$$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}})$ (65)
@@
¯()=¯$\overline{\mathbbm{R}}(\mathcal{M})=\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}}).$ (66)
@@
$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{M}^{\prime})\leq\overline{\mathbbm{R}}(\mathcal{M})+\overline{\mathbbm{R}}(\mathcal{M}^{\prime})$ for any =$\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H}_{A})\}$ and =$\mathcal{M}^{\prime}=\{M^{\prime}_{i}\in\mathscr{M}(\mathcal{H}_{B})\}$. In other words, $\overline{\mathbbm{R}}$ is subadditive under tensor product.
@@
as well as used property [D4], in the fifth line we defined $\mathcal{K}_{ABCD}:=\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{C}\otimes\mathcal{H}_{D}$ and also have used the assumptions [A1] and [A4] (and therefore, $\tilde{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{C}\otimes\mathcal{H}_{D}}$ is a free object). Now, note that both $\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})$ and $\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})$ are positive real numbers and therefore, ¯()+¯$\overline{\mathbbm{R}}(\mathcal{M})+\overline{\mathbbm{R}}(\mathcal{M}^{\prime})=\inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+\inf_{\mathcal{H}_{D}}\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})=\inf_{\mathcal{H}_{C}\otimes\mathcal{H}_{D}}[\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})]$. Then we can write,
@@
$\hskip 56.9055pt\geq\mathbbm{K}(\mathcal{M}\otimes\mathcal{M}^{\prime},\mathcal{H}_{C}\otimes\mathcal{H}_{D})\leavevmode\nobreak\ \forall\mathcal{H}_{C},\mathcal{H}_{D}$
@@
$or,$ $\leavevmode\nobreak\ \inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+\inf_{\mathcal{H}_{D}}\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})$
@@
$or,$$\leavevmode\nobreak\ \overline{\mathbbm{R}}(\mathcal{M})+\overline{\mathbbm{R}}(\mathcal{M}^{\prime})\geq\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{M}^{\prime}).$ (71)
@@
For a convex resource theory, ¯$\overline{\mathbbm{R}}(p\mathcal{M}+(1-p)\mathcal{M}^{\prime})\leq p\overline{\mathbbm{R}}(\mathcal{M})+(1-p)\overline{\mathbbm{R}}(\mathcal{M}^{\prime})$ for any =$\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H}_{A})\}$ and =$\mathcal{M}^{\prime}=\{M^{\prime}_{i}\in\mathscr{M}(\mathcal{H}_{A})\}$ and any $0\leq p\leq 1$.
@@
Again let ′′$\mathcal{M}^{\prime\prime}:=p\mathcal{M}+(1-p)\mathcal{M}^{\prime}$. Then for an arbitrary $\mathcal{H}_{C}$
@@
where in the third line, we have used the Property [D1], in the fourth line, we used the obvious fact that ′′^$\hat{\mathcal{M}^{\prime\prime}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}}=p\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}}+(1-p)\hat{\mathcal{M}}^{\prime}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}}$, and we have used the fact that as the resource theory is convex, free objects form a convex set and therefore, $p\bar{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}}+(1-p)\bar{\mathcal{N}}^{\prime}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}}$ is free object.
@@
$or,\leavevmode\nobreak$ $\inf_{\mathcal{H}_{C}}[p\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+(1-p)\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{C})]$
@@
$or,\leavevmode\nobreak$ $\inf_{\mathcal{H}_{C}}p\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+\inf_{\mathcal{H}_{C}}(1-p)\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{C})$
@@
$or,\leavevmode\nobreak$$\overline{\mathbbm{R}}(\mathcal{M})+\overline{\mathbbm{R}}(\mathcal{M}^{\prime})$
¯$\hskip 56.9055pt\geq\overline{\mathbbm{R}}(p\mathcal{M}+(1-p)\mathcal{M}^{\prime}).$ (75)
Where in the second line we have used the fact that ¯()+¯$\overline{\mathbbm{R}}(\mathcal{M})+\overline{\mathbbm{R}}(\mathcal{M}^{\prime})=\inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+\inf_{\mathcal{H}_{D}}\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})=\inf_{\mathcal{H}_{C}\otimes\mathcal{H}_{D}}[\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})]$ as both $\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})$ and $\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})$ are positive real numbers. ∎
@@
If marginalization is a free transformation (i.e., if assumption [A6] holds) for a given resource theory, then, $\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{M}^{\prime})\geq\overline{\mathbbm{R}}(\mathcal{M})$ for any =$\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H}_{A})\}$ and any $\mathcal{N}=\{N_{i}\in\mathscr{M}(\mathcal{H}_{B})\}\in\mathscr{M}(\mathcal{H}_{A})$.
@@
$\mathbbm{K}(\mathcal{M}\otimes\mathcal{M}^{\prime},\mathcal{H}_{C})\geq\mathbbm{K}(\mathcal{M},\mathcal{H}_{B}\otimes\mathcal{H}_{C})\leavevmode\nobreak\ \forall\mathcal{H}_{C}$
$or,\leavevmode\nobreak$ $\inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M}\otimes\mathcal{M}^{\prime},\mathcal{H}_{C})\geq\inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{B}\otimes\mathcal{H}_{C})$
$or,\leavevmode\nobreak$$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{M}^{\prime})\geq\inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{B}\otimes\mathcal{H}_{C})$
$or,\leavevmode\nobreak$$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{M}^{\prime})\geq\inf_{\mathcal{H}_{B}\otimes\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{B}\otimes\mathcal{H}_{C})$
$or,\leavevmode\nobreak$$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{M}^{\prime})\geq\overline{\mathbbm{R}}(\mathcal{M}).$ (78)
@@
If marginalization is a free transformation (i.e., if assumption [A6] holds) for a given resource theory, then, we have ¯$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{N})=\overline{\mathbbm{R}}(\mathcal{M})$ for any =$\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H}_{A})\}$ and any $\mathcal{N}=\{N_{i}\in\mathscr{M}(\mathcal{H}_{B})\}\in\mathcal{F}_{\mathcal{H}_{B}}$.
@@
As $\mathcal{N}\in\mathcal{F}_{\mathcal{H}_{B}}$, ¯$\overline{\mathbbm{R}}(\mathcal{N})=0$. Therefore, from Proposition 3, we get,
¯$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{N})\leq\overline{\mathbbm{R}}(\mathcal{M}).$ (79)
@@
¯$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{N})\geq\overline{\mathbbm{R}}(\mathcal{M}).$ (80)
@@
¯$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{N})=\overline{\mathbbm{R}}(\mathcal{M}).$ (81)
@@
If $\mathbbm{D}=\overline{\mathcal{D}}$, then for any set of measurement $\mathcal{M}\subset\mathscr{M}(\mathcal{H}_{A})$, we have
$\overline{\mathbbm{R}}(\mathcal{M})\leq$ $\min\left\{\frac{2\mathscr{R}(\mathcal{M})}{1+\mathscr{R}(\mathcal{M})},\frac{2\mathscr{W}(\mathcal{M})}{1+\mathscr{W}(\mathcal{M})}.\right\}$ (82)
@@
We have $\mathbbm{D}=\overline{\mathcal{D}}$. Suppose in Eq. (23), the minimum occurs for ~==$\tilde{\mathcal{M}}=\mathcal{M}^{*}=\{M_{i}^{*}\}$ and $\mathscr{R}(\mathcal{M})=r^{*}$. Then
$N_{i}=\frac{M_{i}}{1+r^{*}}+\frac{r^{*}M^{*}_{i}}{1+r^{*}}\leavevmode\nobreak\ \forall i$
$or,\leavevmode\nobreak\ N_{i}\otimes\mathcal{T}_{\mathcal{B}}=\frac{M_{i}\otimes\mathcal{T}_{\mathcal{B}}}{1+r^{*}}+\frac{r^{*}\tilde{M}_{i}\otimes\mathcal{T}_{\mathcal{B}}}{1+r^{*}}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ (\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}=\frac{(\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}{1+r^{*}}+\frac{r^{*}(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}{1+r^{*}}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ (\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$
$\hskip 56.9055pt=\frac{r^{*}[(\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}]}{1+r^{*}}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ \mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\hskip 56.9055pt=\frac{r^{*}}{1+r^{*}}\mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ \mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\hskip 56.9055pt\leq\frac{2r^{*}}{1+r^{*}}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$ (83)
In the last inequality, we have used the fact that $\mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})\leq 2$ (see Remark 2). Let
$\overline{\mathcal{D}}(\hat{M}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\hat{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})=\mathcal{D}_{\Diamond}((\hat{M}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}),$ (84)
@@
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})=$ $\min_{\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}\overline{\mathcal{D}}(\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}},\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}).$ (85)
@@
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})$ $\leq\overline{\mathcal{D}}(\hat{M}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\hat{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$=\mathcal{D}_{\Diamond}((\hat{M}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\frac{2r^{*}}{1+r^{*}}.$ (86)

### v5

@@ III.2 Distance-based measures for measurement-based resources
For the simplicity of notation in the following arguments let us denote $\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}}$ as $\mathcal{M}^{\prime}$. Then, we have $\hat{\mathcal{M}^{\prime}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{C}}=\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{C}}$ and therefore we can write the following,
@@
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})\geq$ $\mathbbm{K}(\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}},\mathcal{H}_{C})\penalty 10000\ \forall\mathcal{H}_{C}$
$or,\penalty 10000\ \inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})\geq$ $\inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}},\mathcal{H}_{C})$
$or,\penalty 10000\ \overline{\mathbbm{R}}(\mathcal{M})\geq$ $\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}})$ (65)
@@
$\overline{\mathbbm{R}}(\mathcal{M})=\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{T}_{\mathcal{H}_{B}}).$ (66)
@@
[…]
@@
as well as used property [D4], in the fifth line we defined $\mathcal{K}_{ABCD}:=\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{C}\otimes\mathcal{H}_{D}$ and also have used the assumptions [A1] and [A4] (and therefore, $\tilde{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{C}\otimes\mathcal{H}_{D}}$ is a free object). Now, note that both $\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})$ and $\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})$ are positive real numbers and therefore, $\overline{\mathbbm{R}}(\mathcal{M})+\overline{\mathbbm{R}}(\mathcal{M}^{\prime})=\inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+\inf_{\mathcal{H}_{D}}\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})=\inf_{\mathcal{H}_{C}\otimes\mathcal{H}_{D}}[\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})]$. Then we can write,
@@
$\hskip 56.9055pt\geq\mathbbm{K}(\mathcal{M}\otimes\mathcal{M}^{\prime},\mathcal{H}_{C}\otimes\mathcal{H}_{D})\penalty 10000\ \forall\mathcal{H}_{C},\mathcal{H}_{D}$
@@
$or,$ $\penalty 10000\ \inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+\inf_{\mathcal{H}_{D}}\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})$
@@
[…]
@@
For a convex resource theory, $\overline{\mathbbm{R}}(p\mathcal{M}+(1-p)\mathcal{M}^{\prime})\leq p\overline{\mathbbm{R}}(\mathcal{M})+(1-p)\overline{\mathbbm{R}}(\mathcal{M}^{\prime})$ for any $\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H}_{A})\}$ and $\mathcal{M}^{\prime}=\{M^{\prime}_{i}\in\mathscr{M}(\mathcal{H}_{A})\}$ and any $0\leq p\leq 1$.
@@
Again let $\mathcal{M}^{\prime\prime}:=p\mathcal{M}+(1-p)\mathcal{M}^{\prime}$. Then for an arbitrary $\mathcal{H}_{C}$
@@
where in the third line, we have used the Property [D1], in the fourth line, we used the obvious fact that $\hat{\mathcal{M}^{\prime\prime}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}}=p\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}}+(1-p)\hat{\mathcal{M}}^{\prime}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}}$, and we have used the fact that as the resource theory is convex, free objects form a convex set and therefore, $p\bar{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}}+(1-p)\bar{\mathcal{N}}^{\prime}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}}$ is free object.
@@
$or,\penalty 10000\$$\inf_{\mathcal{H}_{C}}[p\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+(1-p)\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{C})]$
@@
$or,\penalty 10000\$$\inf_{\mathcal{H}_{C}}p\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+\inf_{\mathcal{H}_{C}}(1-p)\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{C})$
@@
$or,\penalty 10000\$$p\overline{\mathbbm{R}}(\mathcal{M})+(1-p)\overline{\mathbbm{R}}(\mathcal{M}^{\prime})$
$\hskip 56.9055pt\geq\overline{\mathbbm{R}}(p\mathcal{M}+(1-p)\mathcal{M}^{\prime}).$ (75)
Where in the second line we have used the fact that $\overline{\mathbbm{R}}(\mathcal{M})+\overline{\mathbbm{R}}(\mathcal{M}^{\prime})=\inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+\inf_{\mathcal{H}_{D}}\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})=\inf_{\mathcal{H}_{C}\otimes\mathcal{H}_{D}}[\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})+\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})]$ as both $\mathbbm{K}(\mathcal{M},\mathcal{H}_{C})$ and $\mathbbm{K}(\mathcal{M}^{\prime},\mathcal{H}_{D})$ are positive real numbers. ∎
@@
[…]
@@
$\mathbbm{K}(\mathcal{M}\otimes\mathcal{M}^{\prime},\mathcal{H}_{C})\geq\mathbbm{K}(\mathcal{M},\mathcal{H}_{B}\otimes\mathcal{H}_{C})\penalty 10000\ \forall\mathcal{H}_{C}$
$or,\penalty 10000\$$\inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M}\otimes\mathcal{M}^{\prime},\mathcal{H}_{C})\geq\inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{B}\otimes\mathcal{H}_{C})$
$or,\penalty 10000\$$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{M}^{\prime})\geq\inf_{\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{B}\otimes\mathcal{H}_{C})$
$or,\penalty 10000\$$\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{M}^{\prime})\geq\inf_{\mathcal{H}_{B}\otimes\mathcal{H}_{C}}\mathbbm{K}(\mathcal{M},\mathcal{H}_{B}\otimes\mathcal{H}_{C})$
[…]
@@
If marginalization is a free transformation (i.e., if assumption [A6] holds) for a given resource theory, then, we have $\overline{\mathbbm{R}}(\mathcal{M}\otimes\mathcal{N})=\overline{\mathbbm{R}}(\mathcal{M})$ for any $\mathcal{M}=\{M_{i}\in\mathscr{M}(\mathcal{H}_{A})\}$ and any $\mathcal{N}=\{N_{i}\in\mathscr{M}(\mathcal{H}_{B})\}\in\mathcal{F}_{\mathcal{H}_{B}}$.
@@
As $\mathcal{N}\in\mathcal{F}_{\mathcal{H}_{B}}$, $\overline{\mathbbm{R}}(\mathcal{N})=0$. Therefore, from Proposition 3, we get,
[…]
@@
[…]
@@
[…]
@@
If $\mathbbm{D}=\widetilde{\mathcal{D}}$, then for any set of measurements $\mathcal{M}\subset\mathscr{M}(\mathcal{H}_{A})$, we have
@@
We have $\mathbbm{D}=\widetilde{\mathcal{D}}$. Suppose in Eq. (23), the minimum occurs for $\tilde{\mathcal{M}}=\mathcal{M}^{*}=\{M_{i}^{*}\}$ and $\mathscr{R}(\mathcal{M})=r^{*}$. Then
$N_{i}=\frac{M_{i}}{1+r^{*}}+\frac{r^{*}M^{*}_{i}}{1+r^{*}}\penalty 10000\ \forall i$
$or,\penalty 10000\ N_{i}\otimes\mathcal{T}_{\mathcal{B}}=\frac{M_{i}\otimes\mathcal{T}_{\mathcal{B}}}{1+r^{*}}+\frac{r^{*}\tilde{M}_{i}\otimes\mathcal{T}_{\mathcal{B}}}{1+r^{*}}\penalty 10000\ \forall i,\mathcal{H}_{B}$
$or,\penalty 10000\ (\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}=\frac{(\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}{1+r^{*}}+\frac{r^{*}(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}{1+r^{*}}\penalty 10000\ \forall i,\mathcal{H}_{B}$
$or,\penalty 10000\ (\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$
$\hskip 56.9055pt=\frac{r^{*}[(\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}]}{1+r^{*}}\penalty 10000\ \forall i,\mathcal{H}_{B}$
$or,\penalty 10000\ \mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\hskip 56.9055pt=\frac{r^{*}}{1+r^{*}}\mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})\penalty 10000\ \forall i,\mathcal{H}_{B}$
$or,\penalty 10000\ \mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\hskip 56.9055pt\leq\frac{2r^{*}}{1+r^{*}}\penalty 10000\ \forall i,\mathcal{H}_{B}$ (83)
In the last inequality, we have used the fact that $\mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})\leq 2$ (see Remark 2). Let
$\widetilde{\mathcal{D}}(\hat{M}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\hat{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})=\mathcal{D}_{\Diamond}((\hat{M}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}),$ (84)
@@
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})=$ $\min_{\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}\widetilde{\mathcal{D}}(\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}},\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}).$ (85)
@@
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})$ $\leq\widetilde{\mathcal{D}}(\hat{M}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\hat{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$=\mathcal{D}_{\Diamond}((\hat{M}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\leq\frac{2r^{*}}{1+r^{*}}.$ (86)